OpenStudy (anonymous):

An observer notices that an aeroplane flies directly overhead. Two minutes later the aeroplane is at an angle of elevation of 25 degrees. If the speed of the aeroplane is 350 m/s, find the altitude of the aeroplane.

OpenStudy (dumbcow):

 d is distance traveled d = v*t d = 350*120 using trig we can solve for h tan(65) = d/h --> h = d/tan(65) hope that helps :)